Using any method you prefer, find the IVP solution of the given probler y + 4y = 0; y(0) = - 10, y(0) = 36 The solution i
Answer #1

Solutions ys 4y=0 40)2-10 & 9) =36 » y + 4y =0 - 6%AD)y=0 fuciliary equation mam 20 m(m + 4) 20 mzo and m+420 m =-4 m2 0,-Y(0)2 – 46 e 4(0) Now, if equation is the form of it CF = cemat Gemet (-4) + e Colt ce 2 Ca e C.F. 2 C, + Get and also P. I =36 = 4 x L Ca 2 - 36 4 -9 Ga-9 now put the value of ca in eqı (b) G+q=-10 9-92-10 C,= 10+9 Ca E-l Now put the value of c, and